OpenStudy (anonymous):

What is the justification of 3x+6=14-x

OpenStudy (anonymous):

HI Welcome to Openstudy! first thing you have to do is to "collect like terms" \(3x+6=14-x\\3x+x=14-6\) then simplify to solve for x :) \(3x+x=14-6\\4x=8\\x=?\)

OpenStudy (anonymous):

what wil you get?